# The Prevalence of Neural Collapse in Neural Multivariate Regression

### v1

@@ 1 Introduction
Regression in deep learning is arguably equally important as classification, as it serves for numerous applications across diverse domains. In imitation learning for autonomous driving, regression is employed to predict continuous control actions (such as steering angles and speed) based on observed human driver behavior. Similarly, regression is used in robotics, where the regression model is trained to imitate expert demonstrations. In the financial sector, regression models are extensively used for predictive analytics, such as forecasting stock prices, estimating risk, and predicting market trends. Meteorology also heavily relies on regression models to forecast weather conditions. These models take high-dimensional inputs from various sensors and satellites to predict multiple continuous variables such as temperature, humidity, and wind speed. Moreover, many reinforcement learning algorithms include critical regression components, where regression is employed to predict value functions with the targets being Monte Carlo or bootstrapped returns.
@@
To the best of our knowledge, this is the first empirical and theoretical study of neural collapse in the context of regression. By demonstrating the prevalence of neural collapse in regression tasks, we reveal that deep learning systems might inherently simplify their internal representations, irrespective of the specific nature of the task, whether it be classification or regression. This simplification can lead to more explainable models and potentially more efficient training processes.
@@ 3 Prevalence of neural regression collapse
We train the DDN using gradient descent to minimize the regularized L2 loss:
@@ 3.1 Definition of neural regression collapse
Also denote ${\mathbf{H}}:=[{\mathbf{h}}_{1}\cdots{\mathbf{h}}_{M}$], where ${\mathbf{h}}_{i}$ is the feature vector associated with input ${\mathbf{x}}_{i}$, that is, ${\mathbf{h}}_{i}:={\mathbf{h}}_{\mathbf{\theta}}({\mathbf{x}}_{i})$. Of course, ${\mathbf{W}}$, ${\mathbf{H}}$, and ${\mathbf{b}}$ are changing throughout the course of training. For any $p\times q$ matrix ${\mathbf{C}}$ and any $p$-dimensional vector ${\mathbf{v}}$, let $proj({\mathbf{v}}|{\mathbf{C}})$ denote the projection of ${\mathbf{v}}$ onto the subspace spanned by the columns of ${\mathbf{C}}$. Let ${\mathbf{H}}_{PCA_{n}}$ be the $d\times n$ matrix with the columns consisting of the $n$ principal components of ${\mathbf{H}}$.
@@
  • NRC1 = $\frac{1}{M}\sum_{i=1}^{M}||{\mathbf{h}}_{i}-proj({\mathbf{h}}_{i}|{\mathbf{H}}_{PCA_{n}})||_{2}^{2}\to 0$.
  • NRC2 = $\frac{1}{M}\sum_{i=1}^{M}||{\mathbf{h}}_{i}-proj({\mathbf{h}}_{i}|{\mathbf{W}}^{T})||_{2}^{2}\to 0$.
@@ 3.2 Experimental validation of neural regression collapse
  • The CARLA dataset originates from the CARLA Simulator, an open-source project designed to support the development of autonomous driving systems. We utilize a dataset Codevilla et al. (2018) sourced from expert-driven offline simulations. During these simulations, images (${\mathbf{x}}_{i}$’s) from cameras mounted on the virtual vehicle and corresponding expert driver actions as targets (${\mathbf{y}}_{i}$’s) are recorded as human drives in the simulated environment. We consider two dataset versions: a 2D version with steering angle and speed, and a 1D version with only the steering angle.
@@
All experimental results are averaged over 3 random seeds and variance is displayed by a shaded area. The choices of weight decay employed during training varied depending on the dataset. Also, the number of epochs required for training depends on both the dataset and the degree of weight decay. In particular, we used a large number of epochs when using very small weight decay values. Appendix A provides the full experimental setup.
Empirical Results. Figure 2 presents the experimental results for the six datasets mentioned above. The results show that the training and testing errors decrease as training progresses, as expected. The converging coefficient of determination ($R^{2}$) also indicates that model performance becomes stable. Most importantly, the figure confirms the presence of NRC1-NRC3 across all six datasets. This indicates that neural collapse is not only prevalent in classification but also often occurs in multivariate regression. In Appendix A.3, we also experimentally analyze the explained variance ratio of principal components to further verify the collapse to the subspace spanned by the first $n$ components. Our definition of NRC3 involves finding a scaling factor $\gamma$ for which the property holds. Figure 3 illustrates the values of NRC3 as a function of $\gamma$ for ${\mathbf{W}}$ obtained after training. We observe that each dataset exhibits a unique minimum value of $\gamma$. More details about computing NRC3 can be found in Appendix A.4.
@@ 4.3 General $n$-dimensional multivariate case
Suppose $0<c<\lambda_{\text{min}}$. Then the global minima given by (3) have the following properties: (i) All of the $d$-dimensional feature vectors ${\mathbf{h}}_{i}$, $i=1,\ldots,M$, lie in the $n$-dimensional subspace spanned by the $n$ rows of ${\mathbf{W}}$. (ii) ${\mathbf{W}}{\mathbf{W}}^{T}=\sqrt{\frac{\lambda_{{\mathbf{H}}}}{\lambda_{{\mathbf{W}}}}}\left[{\mathbf{\Sigma}}^{1/2}-\sqrt{c}{\mathbf{I}}_{n}\right]$, (iii) $\lambda_{{\mathbf{H}}}||{\mathbf{H}}||_{F}^{2}=M\lambda_{{\mathbf{W}}}||{\mathbf{W}}||_{F}^{2}$, (iv) $\mathcal{L}({\mathbf{H}},{\mathbf{W}},{\mathbf{b}})=\frac{nc}{2}+\lambda_{{\mathbf{W}}}||{\mathbf{A}}^{1/2}||_{F}^{2}$, (v) ${\mathbf{W}}{\mathbf{H}}+{\mathbf{b}}\mathbf{1}_{M}^{T}-{\mathbf{Y}}=-\sqrt{c}[{\mathbf{\Sigma}}^{1/2}]^{-1}({\mathbf{Y}}-{\bar{\mathbf{Y}}})$.
@@ A.2 CARLA and UTKface
The Carla dataset is collected by recording surroundings via automotive cameras, while a human driver operates a vehicle in a simulative urban environment [Codevilla et al., 2018]. The recorded images are states ${\mathbf{x}}_{i}$ of the vehicle and the expert control from the driver, which includes steering angles and speed, serves as actions ${\mathbf{y}}_{i}\in[-1,1]\times[0,85]$ in the dataset. A well-trained model on this dataset is expected to drive the vehicle safely in the virtual environment. The UTKface dataset consists of full-face photographs ${\mathbf{x}}_{i}$ of humans whose ages range from 1 to 116 [Zhang et al., 2017]. The goal of this dataset is to accurately predict the age ${\mathbf{y}}_{i}$ of the person in each photo.
@@
Number of compute workers 1
@@
A.3 Analysis on NRC1 definition
We provide additional experimental justification for choosing the dimension of collapsed subspace in the NRC1 definition to be equal to $n$, the dimension of the targets. Recall that NRC1 is defined as:
[…]
where ${\mathbf{H}}_{PCA_{n}}$ is the $d\times n$ matrix with the columns consisting of the $n$ principal components of ${\mathbf{H}}$. In Figure 6, we investigate the explained variance ratio (EVR) of the first 5 principal components of ${\mathbf{H}}$ during the training process. For all datasets, there is significant variance for all of the first $n$ components after a short period of training; for other components, there is very low or even no variance. This indicates that a perfect collapse occurs in the subspace spanned by the first $n$ principal components.
Figure 6: Explained Variance Ratio for the first 5 principal components (PC).
A.4 Computing NRC3
@@
A.5 Results for small weight decay

### v2

@@ 1 Introduction
Regression in deep learning is arguably equally important as classification, as it serves for numerous applications across diverse domains. In imitation learning for autonomous driving, regression is employed to predict continuous control actions (such as speed and steering angles) based on observed human driver behavior. Similarly, regression is used in robotics, where the regression model is trained to imitate expert demonstrations. In the financial sector, regression models are extensively used for predictive analytics, such as forecasting stock prices, estimating risk, and predicting market trends. Meteorology also heavily relies on regression models to forecast weather conditions. These models take high-dimensional inputs from various sensors and satellites to predict multiple continuous variables such as temperature, humidity, and wind speed. Moreover, many reinforcement learning algorithms include critical regression components, where regression is employed to predict value functions with the targets being Monte Carlo or bootstrapped returns.
@@
To the best of our knowledge, this is the first empirical and theoretical study of neural collapse in the context of regression. By demonstrating the prevalence of neural collapse in regression tasks, we reveal that deep learning systems might inherently simplify their internal representations, irrespective of the specific nature of the task, whether it be classification or regression.
@@ 3 Prevalence of neural regression collapse
We train the DNN using gradient descent to minimize the regularized L2 loss:
@@ 3.1 Definition of neural regression collapse
Denote ${\mathbf{H}}:=[{\mathbf{h}}_{1}\cdots{\mathbf{h}}_{M}$], where ${\mathbf{h}}_{i}$ is the feature vector associated with input ${\mathbf{x}}_{i}$, that is, ${\mathbf{h}}_{i}:={\mathbf{h}}_{\mathbf{\theta}}({\mathbf{x}}_{i})$. Further denote the normalized feature vector $\widetilde{{\mathbf{h}}}_{i}:={\mathbf{h}}_{i}\cdot||{\mathbf{h}}_{i}||^{-1}$. Of course, ${\mathbf{W}}$, ${\mathbf{H}}$, and ${\mathbf{b}}$ are changing throughout the course of training. For any $p\times q$ matrix ${\mathbf{C}}$ and any $p$-dimensional vector ${\mathbf{v}}$, let $proj({\mathbf{v}}|{\mathbf{C}})$ denote the projection of ${\mathbf{v}}$ onto the subspace spanned by the columns of ${\mathbf{C}}$. Let ${\mathbf{H}}_{PCA_{n}}$ be the $d\times n$ matrix with the columns consisting of the $n$ principal components of ${\mathbf{H}}$.
@@
  • NRC1 = $\frac{1}{M}\sum_{i=1}^{M}\left|\left|\widetilde{{\mathbf{h}}}_{i}-proj(\widetilde{{\mathbf{h}}}_{i}|{\mathbf{H}}_{PCA_{n}})\right|\right|_{2}^{2}\to 0$.
  • NRC2 = $\frac{1}{M}\sum_{i=1}^{M}\left|\left|\widetilde{{\mathbf{h}}}_{i}-proj(\widetilde{{\mathbf{h}}}_{i}|{\mathbf{W}}^{T})\right|\right|_{2}^{2}\to 0$.
@@ 3.2 Experimental validation of neural regression collapse
  • The CARLA dataset originates from the CARLA Simulator, an open-source project designed to support the development of autonomous driving systems. We utilize a dataset Codevilla et al. (2018) sourced from expert-driven offline simulations. During these simulations, images (${\mathbf{x}}_{i}$’s) from cameras mounted on the virtual vehicle and corresponding expert driver actions as targets (${\mathbf{y}}_{i}$’s) are recorded as human drives in the simulated environment. We consider two dataset versions: a 2D version with speed and steering angle, and a 1D version with only the speed.
@@
All experimental results are averaged over at least 2 random seeds and variance is displayed by a shaded area. The choices of weight decay employed during training varied depending on the dataset. Also, the number of epochs required for training depends on both the dataset and the degree of weight decay. In particular, we used a large number of epochs when using very small weight decay values. appendix A provides the full experimental setup.
Empirical Results. Figure 2 presents the experimental results for the six datasets mentioned above. The results show that the training and testing errors decrease as training progresses, as expected. The converging coefficient of determination ($R^{2}$) also indicates that model performance becomes stable. Most importantly, the figure confirms the presence of NRC1-NRC3 across all six datasets. This indicates that neural collapse is not only prevalent in classification but also often occurs in multivariate regression.
@@
Figure 3: Explained Variance Ratio (EVR) for the first 5 principal components (PC).
We also experimentally analyze the explained variance ratio (EVR) of principal components to further verify the collapse to the subspace spanned by the first $n$ components. In Figure 3, we investigate the EVR of the first 5 principal components of ${\mathbf{H}}$ during the training process. For all datasets, there is significant variance for all of the first $n$ components after a short period of training; for other components, there is very low or even no variance. This also supports that a perfect collapse occurs in the subspace spanned by the first $n$ principal components.
Our definition of NRC3 involves finding a scaling factor $\gamma$ for which the property holds. Figure 4 illustrates the values of NRC3 as a function of $\gamma$ for ${\mathbf{W}}$ obtained after training. We observe that each dataset exhibits a unique minimum value of $\gamma$. More details about computing NRC3 can be found in appendix A.3.
@@ 4.3 General $n$-dimensional multivariate case
Suppose $0<c<\lambda_{\text{min}}$. Then the global minima given by (3) have the following properties:
(i) All of the $d$-dimensional feature vectors ${\mathbf{h}}_{i}$, $i=1,\ldots,M$, lie in the $n$-dimensional subspace spanned by the $n$ rows of ${\mathbf{W}}$. (ii) ${\mathbf{W}}{\mathbf{W}}^{T}=\sqrt{\frac{\lambda_{{\mathbf{H}}}}{\lambda_{{\mathbf{W}}}}}\left[{\mathbf{\Sigma}}^{1/2}-\sqrt{c}{\mathbf{I}}_{n}\right]$, (iii) $\lambda_{{\mathbf{H}}}||{\mathbf{H}}||_{F}^{2}=M\lambda_{{\mathbf{W}}}||{\mathbf{W}}||_{F}^{2}$, (iv) $\mathcal{L}({\mathbf{H}},{\mathbf{W}},{\mathbf{b}})=nc/2+\sqrt{c}||{\mathbf{A}}^{1/2}||_{F}^{2}$, (v) ${\mathbf{W}}{\mathbf{H}}+{\mathbf{b}}\mathbf{1}_{M}^{T}-{\mathbf{Y}}=-\sqrt{c}[{\mathbf{\Sigma}}^{1/2}]^{-1}({\mathbf{Y}}-{\bar{\mathbf{Y}}})$.
@@ 5 Conclusion
This work is submitted in part by the NYU Abu Dhabi Center for Artificial Intelligence and Robotics, funded by Tamkeen under the Research Institute Award CG010.
@@ A.2 CARLA and UTKface
The Carla dataset is collected by recording surroundings via automotive cameras, while a human driver operates a vehicle in a simulative urban environment [Codevilla et al., 2018]. The recorded images are states ${\mathbf{x}}_{i}$ of the vehicle and the expert control from the driver, which includes speed and steering angles, serves as actions ${\mathbf{y}}_{i}\in[0,85]\times[-1,1]$ in the dataset. A well-trained model on this dataset is expected to drive the vehicle safely in the virtual environment. The UTKface dataset consists of full-face photographs ${\mathbf{x}}_{i}$ of humans whose ages range from 1 to 116 [Zhang et al., 2017]. The goal of this dataset is to accurately predict the age ${\mathbf{y}}_{i}$ of the person in each photo.
@@
Number of compute workers 8
@@
A.3 Computing NRC3
@@
A.4 Results for small weight decay
@@ Appendix F Proof of the uniqueness of $\gamma$
NeurIPS Paper Checklist
  1. 1.
    Claims
  2. Question: Do the main claims made in the abstract and introduction accurately reflect the paper’s contributions and scope?
  3. Answer: [Yes]
  4. Justification: In the abstract, we claim to empirically demonstrate the phenomenon of Neural Regression Collapse as discussed in Section 3.2, and then we provide a theoretical explanation from the perspective of the Unconstrained Feature Model, see Section 4 and Appendix B,C,D. Thus, we show that the phenomena of neural collapse could be a universal behavior in deep learning both empirically and theoretically.
  5. Guidelines:
    • The answer NA means that the abstract and introduction do not include the claims made in the paper.
    • The abstract and/or introduction should clearly state the claims made, including the contributions made in the paper and important assumptions and limitations. A No or NA answer to this question will not be perceived well by the reviewers.
    • The claims made should match theoretical and experimental results, and reflect how much the results can be expected to generalize to other settings.
    • It is fine to include aspirational goals as motivation as long as it is clear that these goals are not attained by the paper.
  6. 2.
    Limitations
  7. Question: Does the paper discuss the limitations of the work performed by the authors?
  8. Answer: [Yes]
  9. Justification: In Section 5, we point out that our explanation of neural regression collapse doesn’t have implications for model generalization.
  10. Guidelines:
    • The answer NA means that the paper has no limitation while the answer No means that the paper has limitations, but those are not discussed in the paper.
    • The authors are encouraged to create a separate "Limitations" section in their paper.
    • The paper should point out any strong assumptions and how robust the results are to violations of these assumptions (e.g.,, independence assumptions, noiseless settings, model well-specification, asymptotic approximations only holding locally). The authors should reflect on how these assumptions might be violated in practice and what the implications would be.
    • The authors should reflect on the scope of the claims made, e.g., if the approach was only tested on a few datasets or with a few runs. In general, empirical results often depend on implicit assumptions, which should be articulated.
    • The authors should reflect on the factors that influence the performance of the approach. For example, a facial recognition algorithm may perform poorly when the image resolution is low or images are taken in low lighting. Or a speech-to-text system might not be used reliably to provide closed captions for online lectures because it fails to handle technical jargon.
    • The authors should discuss the computational efficiency of the proposed algorithms and how they scale with dataset size.
    • If applicable, the authors should discuss possible limitations of their approach to address problems of privacy and fairness.
    • While the authors might fear that complete honesty about limitations might be used by reviewers as grounds for rejection, a worse outcome might be that reviewers discover limitations that aren’t acknowledged in the paper. The authors should use their best judgment and recognize that individual actions in favor of transparency play an important role in developing norms that preserve the integrity of the community. Reviewers will be specifically instructed to not penalize honesty concerning limitations.
  11. 3.
    Theory Assumptions and Proofs
  12. Question: For each theoretical result, does the paper provide the full set of assumptions and a complete (and correct) proof?
  13. Answer: [Yes]
  14. Justification: In Section 4 and Appendix B,C,D, E, we provide full set of assumptions and proofs.
  15. Guidelines:
    • The answer NA means that the paper does not include theoretical results.
    • All the theorems, formulas, and proofs in the paper should be numbered and cross-referenced.
    • All assumptions should be clearly stated or referenced in the statement of any theorems.
    • The proofs can either appear in the main paper or the supplemental material, but if they appear in the supplemental material, the authors are encouraged to provide a short proof sketch to provide intuition.
    • Inversely, any informal proof provided in the core of the paper should be complemented by formal proofs provided in the appendix or supplemental material.
    • Theorems and Lemmas that the proof relies upon should be properly referenced.
  16. 4.
    Experimental Result Reproducibility
  17. Question: Does the paper fully disclose all the information needed to reproduce the main experimental results of the paper to the extent that it affects the main claims and/or conclusions of the paper (regardless of whether the code and data are provided or not)?
  18. Answer: [Yes]
  19. Justification: In Sections 3.2 and Appendix A, we provide a comprehensive description of the experimental setup, including detailed information on the datasets used, model architectures, and hyperparameter settings. This detailed disclosure ensures that other researchers can reliably reproduce the experimental results and validate the claims made in the paper.
  20. Guidelines:
    • The answer NA means that the paper does not include experiments.
    • If the paper includes experiments, a No answer to this question will not be perceived well by the reviewers: Making the paper reproducible is important, regardless of whether the code and data are provided or not.
    • If the contribution is a dataset and/or model, the authors should describe the steps taken to make their results reproducible or verifiable.
    • Depending on the contribution, reproducibility can be accomplished in various ways. For example, if the contribution is a novel architecture, describing the architecture fully might suffice, or if the contribution is a specific model and empirical evaluation, it may be necessary to either make it possible for others to replicate the model with the same dataset, or provide access to the model. In general. releasing code and data is often one good way to accomplish this, but reproducibility can also be provided via detailed instructions for how to replicate the results, access to a hosted model (e.g.,, in the case of a large language model), releasing of a model checkpoint, or other means that are appropriate to the research performed.
    • While NeurIPS does not require releasing code, the conference does require all submissions to provide some reasonable avenue for reproducibility, which may depend on the nature of the contribution. For example
      1. (a)
        If the contribution is primarily a new algorithm, the paper should make it clear how to reproduce that algorithm.
      2. (b)
        If the contribution is primarily a new model architecture, the paper should describe the architecture clearly and fully.
      3. (c)
        If the contribution is a new model (e.g.,, a large language model), then there should either be a way to access this model for reproducing the results or a way to reproduce the model (e.g.,, with an open-source dataset or instructions for how to construct the dataset).
      4. (d)
        We recognize that reproducibility may be tricky in some cases, in which case authors are welcome to describe the particular way they provide for reproducibility. In the case of closed-source models, it may be that access to the model is limited in some way (e.g.,, to registered users), but it should be possible for other researchers to have some path to reproducing or verifying the results.
  21. 5.
    Open access to data and code
  22. Question: Does the paper provide open access to the data and code, with sufficient instructions to faithfully reproduce the main experimental results, as described in supplemental material?
  23. Answer: [Yes]
  24. Justification: We upload the code with environment in the supplemental materials. The datasets used are all open-source datasets.
  25. Guidelines:
    • The answer NA means that paper does not include experiments requiring code.
    • Please see the NeurIPS code and data submission guidelines (https://nips.cc/public/guides/CodeSubmissionPolicy) for more details.
    • While we encourage the release of code and data, we understand that this might not be possible, so “No” is an acceptable answer. Papers cannot be rejected simply for not including code, unless this is central to the contribution (e.g.,, for a new open-source benchmark).
    • The instructions should contain the exact command and environment needed to run to reproduce the results. See the NeurIPS code and data submission guidelines (https://nips.cc/public/guides/CodeSubmissionPolicy) for more details.
    • The authors should provide instructions on data access and preparation, including how to access the raw data, preprocessed data, intermediate data, and generated data, etc.
    • The authors should provide scripts to reproduce all experimental results for the new proposed method and baselines. If only a subset of experiments are reproducible, they should state which ones are omitted from the script and why.
    • At submission time, to preserve anonymity, the authors should release anonymized versions (if applicable).
    • Providing as much information as possible in supplemental material (appended to the paper) is recommended, but including URLs to data and code is permitted.
  26. 6.
    Experimental Setting/Details
  27. Question: Does the paper specify all the training and test details (e.g.,, data splits, hyperparameters, how they were chosen, type of optimizer, etc.) necessary to understand the results?
  28. Answer: [Yes]
  29. Justification: In Sections 3.2 and Appendix A, we provide a comprehensive description of the experimental setup, including detailed information on the datasets used, model architectures, and hyperparameter settings. More details can be found in the code in the supplemental materials.
  30. Guidelines:
    • The answer NA means that the paper does not include experiments.
    • The experimental setting should be presented in the core of the paper to a level of detail that is necessary to appreciate the results and make sense of them.
    • The full details can be provided either with the code, in appendix, or as supplemental material.
  31. 7.
    Experiment Statistical Significance
  32. Question: Does the paper report error bars suitably and correctly defined or other appropriate information about the statistical significance of the experiments?
  33. Answer: [No]
  34. Justification: All the results are reported in terms of learning curves, and each figure includes many plots, so error bars are not reported. But we do run the experiments multiple times and observe very similar performance in terms of NRC, testing loss, etc.
  35. Guidelines:
    • The answer NA means that the paper does not include experiments.
    • The authors should answer "Yes" if the results are accompanied by error bars, confidence intervals, or statistical significance tests, at least for the experiments that support the main claims of the paper.
    • The factors of variability that the error bars are capturing should be clearly stated (for example, train/test split, initialization, random drawing of some parameter, or overall run with given experimental conditions).
    • The method for calculating the error bars should be explained (closed form formula, call to a library function, bootstrap, etc.)
    • The assumptions made should be given (e.g.,, Normally distributed errors).
    • It should be clear whether the error bar is the standard deviation or the standard error of the mean.
    • It is OK to report 1-sigma error bars, but one should state it. The authors should preferably report a 2-sigma error bar than state that they have a 96% CI, if the hypothesis of Normality of errors is not verified.
    • For asymmetric distributions, the authors should be careful not to show in tables or figures symmetric error bars that would yield results that are out of range (e.g., negative error rates).
    • If error bars are reported in tables or plots, The authors should explain in the text how they were calculated and reference the corresponding figures or tables in the text.
  36. 8.
    Experiments Compute Resources
  37. Question: For each experiment, does the paper provide sufficient information on the computer resources (type of compute workers, memory, time of execution) needed to reproduce the experiments?
  38. Answer: [Yes]
  39. Justification: In Appendix A, we provide the details for computation resources.
  40. Guidelines:
    • The answer NA means that the paper does not include experiments.
    • The paper should indicate the type of compute workers CPU or GPU, internal cluster, or cloud provider, including relevant memory and storage.
    • The paper should provide the amount of compute required for each of the individual experimental runs as well as estimate the total compute.
    • The paper should disclose whether the full research project required more compute than the experiments reported in the paper (e.g.,, preliminary or failed experiments that didn’t make it into the paper).
  41. 9.
    Code Of Ethics
  42. Question: Does the research conducted in the paper conform, in every respect, with the NeurIPS Code of Ethics https://neurips.cc/public/EthicsGuidelines?
  43. Answer: [Yes]
  44. Justification: We are sure that the research presented in this paper adheres to the NeurIPS Code of Ethics in all respects.
  45. Guidelines:
    • The answer NA means that the authors have not reviewed the NeurIPS Code of Ethics.
    • If the authors answer No, they should explain the special circumstances that require a deviation from the Code of Ethics.
    • The authors should make sure to preserve anonymity (e.g.,, if there is a special consideration due to laws or regulations in their jurisdiction).
  46. 10.
    Broader Impacts
  47. Question: Does the paper discuss both potential positive societal impacts and negative societal impacts of the work performed?
  48. Answer: [N/A]
  49. Justification: This paper mainly focuses on showing and understanding the neural regression collapse phenomena observed in practical neural networks and unconstrained feature models. No potential negative societal impact is expected of this work.
  50. Guidelines:
    • The answer NA means that there is no societal impact of the work performed.
    • If the authors answer NA or No, they should explain why their work has no societal impact or why the paper does not address societal impact.
    • Examples of negative societal impacts include potential malicious or unintended uses (e.g.,, disinformation, generating fake profiles, surveillance), fairness considerations (e.g.,, deployment of technologies that could make decisions that unfairly impact specific groups), privacy considerations, and security considerations.
    • The conference expects that many papers will be foundational research and not tied to particular applications, let alone deployments. However, if there is a direct path to any negative applications, the authors should point it out. For example, it is legitimate to point out that an improvement in the quality of generative models could be used to generate deepfakes for disinformation. On the other hand, it is not needed to point out that a generic algorithm for optimizing neural networks could enable people to train models that generate Deepfakes faster.
    • The authors should consider possible harms that could arise when the technology is being used as intended and functioning correctly, harms that could arise when the technology is being used as intended but gives incorrect results, and harms following from (intentional or unintentional) misuse of the technology.
    • If there are negative societal impacts, the authors could also discuss possible mitigation strategies (e.g.,, gated release of models, providing defenses in addition to attacks, mechanisms for monitoring misuse, mechanisms to monitor how a system learns from feedback over time, improving the efficiency and accessibility of ML).
  51. 11.
    Safeguards
  52. Question: Does the paper describe safeguards that have been put in place for responsible release of data or models that have a high risk for misuse (e.g.,, pretrained language models, image generators, or scraped datasets)?
  53. Answer: [N/A]
  54. Guidelines:
    • The answer NA means that the paper poses no such risks.
    • Released models that have a high risk for misuse or dual-use should be released with necessary safeguards to allow for controlled use of the model, for example by requiring that users adhere to usage guidelines or restrictions to access the model or implementing safety filters.
    • Datasets that have been scraped from the Internet could pose safety risks. The authors should describe how they avoided releasing unsafe images.
    • We recognize that providing effective safeguards is challenging, and many papers do not require this, but we encourage authors to take this into account and make a best faith effort.
  55. 12.
    Licenses for existing assets
  56. Question: Are the creators or original owners of assets (e.g.,, code, data, models), used in the paper, properly credited and are the license and terms of use explicitly mentioned and properly respected?
  57. Answer: [Yes]
  58. Justification: All the owner of assets are properly cited in the reference and main body of our paper.
  59. Guidelines:
    • The answer NA means that the paper does not use existing assets.
    • The authors should cite the original paper that produced the code package or dataset.
    • The authors should state which version of the asset is used and, if possible, include a URL.
    • The name of the license (e.g.,, CC-BY 4.0) should be included for each asset.
    • For scraped data from a particular source (e.g.,, website), the copyright and terms of service of that source should be provided.
    • If assets are released, the license, copyright information, and terms of use in the package should be provided. For popular datasets, paperswithcode.com/datasets has curated licenses for some datasets. Their licensing guide can help determine the license of a dataset.
    • For existing datasets that are re-packaged, both the original license and the license of the derived asset (if it has changed) should be provided.
    • If this information is not available online, the authors are encouraged to reach out to the asset’s creators.
  60. 13.
    New Assets
  61. Question: Are new assets introduced in the paper well documented and is the documentation provided alongside the assets?
  62. Answer: [N/A]
  63. Guidelines:
    • The answer NA means that the paper does not release new assets.
    • Researchers should communicate the details of the dataset/code/model as part of their submissions via structured templates. This includes details about training, license, limitations, etc.
    • The paper should discuss whether and how consent was obtained from people whose asset is used.
    • At submission time, remember to anonymize your assets (if applicable). You can either create an anonymized URL or include an anonymized zip file.
  64. 14.
    Crowdsourcing and Research with Human Subjects
  65. Question: For crowdsourcing experiments and research with human subjects, does the paper include the full text of instructions given to participants and screenshots, if applicable, as well as details about compensation (if any)?
  66. Answer: [N/A]
  67. Guidelines:
    • The answer NA means that the paper does not involve crowdsourcing nor research with human subjects.
    • Including this information in the supplemental material is fine, but if the main contribution of the paper involves human subjects, then as much detail as possible should be included in the main paper.
    • According to the NeurIPS Code of Ethics, workers involved in data collection, curation, or other labor should be paid at least the minimum wage in the country of the data collector.
  68. 15.
    Institutional Review Board (IRB) Approvals or Equivalent for Research with Human Subjects
  69. Question: Does the paper describe potential risks incurred by study participants, whether such risks were disclosed to the subjects, and whether Institutional Review Board (IRB) approvals (or an equivalent approval/review based on the requirements of your country or institution) were obtained?
  70. Answer: [N/A]
  71. Guidelines:
    • The answer NA means that the paper does not involve crowdsourcing nor research with human subjects.
    • Depending on the country in which research is conducted, IRB approval (or equivalent) may be required for any human subjects research. If you obtained IRB approval, you should clearly state this in the paper.
    • We recognize that the procedures for this may vary significantly between institutions and locations, and we expect authors to adhere to the NeurIPS Code of Ethics and the guidelines for their institution.
    • For initial submissions, do not include any information that would break anonymity (if applicable), such as the institution conducting the review.